Implement F(A,B,C)=(A+B+C)(A’+C’)(B+C’) using a 2x1 Multiplexer and using a 4x1 Multiplexer.
Answer #1

F(A,B,C)=(A+B+C)(A'+C')(B+C')

=(A+B+C)(A'+B+C')(A'+B'+C')(A+B+C')(A'+B+C')

=(A+B+C)(A'+B+C')(A'+B'+C')(A+B+C')

=(0+0+0)(1+0+1)(1+1+1)(0+0+1)

=M0 . M5 . M7 . M1

=gif.latex?%5CprodM(0,1,5,7) gif.latex?%5CRightarrowmax terms

F(A,B,C)=gif.latex?%5Csumm(2,3,4,6)   gif.latex?%5CRightarrowmin terms

Do D &X) 戈 n-1) combination) Co 2 3 ㄒㄧ一一て--で

2 3 Au 3